### v1

# LLM4DSR: Leveraing Large Language Model for Denoising Sequential Recommendation

Abstract.
Sequential recommendation systems fundamentally rely on users’ historical interaction sequences, which are often contaminated by noisy interactions. Identifying these noisy interactions accurately without additional information is particularly difficult due to the lack of explicit supervisory signals to denote noise. Large Language Models (LLMs), equipped with extensive open knowledge and semantic reasoning abilities, present a promising avenue to bridge this information gap. However, employing LLMs for denoising in sequential recommendation introduces notable challenges: 1) Direct application of pretrained LLMs may not be competent for the denoising task, frequently generating nonsensical responses; 2) Even after fine-tuning, the reliability of LLM outputs remains questionable, especially given the complexity of the task and th inherent hallucinatory issue of LLMs.
To tackle these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendation using LLMs. We constructed a self-supervised fine-tuning task to activate LLMs’ capabilities to identify noisy items and suggest replacements. Furthermore, we developed an uncertainty estimation module that ensures only high-confidence responses are utilized for sequence corrections. Remarkably, LLM4DSR is model-agnostic, allowing the corrected sequences to be flexibly applied across various recommendation models. Extensive experiments validate the superiority of LLM4DSR over existing methods across three datasets and three recommendation backbones.
1. Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in content comprehension, generation, and semantic reasoning, thereby catalyzing a revolution in artificial intelligence (Achiam et al., 2023). Recently, LLMs have been extensively applied in the field of Recommender Systems (RS) (Wu et al., 2023). Some researchers have directly utilized LLMs by prompting or fine-tuning them to act as specialized recommenders (Liao et al., 2024; Harte et al., 2023; Bao et al., 2023b). Meanwhile, others have employed LLMs as auxiliary tools to enhance traditional recommendation models e.g., serving as encoders for user/item features (Wei et al., 2024), supplementary knowledge bases (Xi et al., 2023), advanced recommendation explainers (Tennenholtz et al., 2023), and conversational agents (Gao et al., 2023).
Given the powerful capabilities of LLMs, this work explores a innovative question: Can LLMs effectively function as denoisers for sequential recommendation? Sequential recommendation critically relies on the accuracy of users’ historical interaction sequences to predict subsequent items. However, in practice, these sequences often contain noise due to various factors — e.g., being attracted by clickbait (Wang et al., 2021b), influenced by item prominent positions (Chen et al., 2023), or from accidental interactions (Lin et al., 2023). Such noise can significantly mislead recommendation models and degrade their performance, thus underscoring the necessity for effectively denoising.
The task of denoising is inherently difficult. A primary issue is the absence of labels for noisy data, resulting in a lack of clear signals regarding the nature of the noise. Recent approaches to denoising in sequential recommendation have either relied on human-designed heuristics (Zhang et al., 2024a, 2022; Sun et al., 2021), which require extensive expertise and often lack precision; or have been integrated with recommendation models guided by the recommendation objective (Zhang et al., 2022; Yuan et al., 2021; Chen et al., 2022; Lin et al., 2023), which may still be heavily influenced by the noisy data, particularly since the labels of the recommendation objective (i.e., the next item) might also be contaminated. Thus, this work explores a new avenue of leveraging LLMs for recommendation denoising, which can potentially compensate for this knowledge gap. LLMs, with their extensive open knowledge and advanced semantic reasoning abilities (Brown et al., 2020), are equipped to identify abnormal interactions from a semantic perspective. For example, if a user’s viewing history primarily consists of comedy films, an occasional interaction with a horror film might be attributable to an accidental click. LLMs can detect such anomalies based on their understanding of item semantics and common knowledge. Moreover, by leveraging the generative capabilities of LLMs, they can not only identify noisy items but also potentially suggest replacements that align more closely with user preferences.
However, employing pre-trained LLMs for sequence denoising also presents significant challenges:
Challenge 1: Direct application of pretrained LLMs may not be competent for denoising tasks. The significant gap between the pre-training objectives of LLMs and the specific requirements of sequence recommendation denoising can make their direct application problematic. Initial tests with a pre-trained Llama3-8B model (Dubey et al., 2024) indicated that a large portion (40%) of the responses were nonsensical. Figure 1 illustrates examples where LLMs typically misidentify all items in a sequence as noise.
Challenge 2: Responses from LLMs might not always be reliable.. Even with fine-tuning, LLMs are susceptible to the notorious phenomenon of hallucination (Ji et al., 2023), where they might incorrectly label non-existent items as noise or suggest non-existent replacements. Moreover, given the complexity of the recommendation denoising task, the accuracy of an LLM-based denoiser still has an upper-bound. Blindly trusting LLMs for corrections may not be advisable.
To address these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendations using LLMs. To overcome the first challenge, we adapt pre-trained LLMs to the denoising task through a self-supervised fine-tuning process. We construct an instruction dataset by replacing a proportion of items in sequences with randomly selected alternatives, which are then used for fine-tuning LLMs to identify noisy items and suggest appropriate replacements. This approach activates the potential denoising ability of LLMs. For the second challenge, we introduce an uncertainty estimation module to assess the reliability of the identified noisy items, ensuring that only high-confidence judgments are used to correct sequences for subsequent recommendations. Remarkably, LLM4DSR is model-agnostic, making the corrected sequences flexibly applicable across various recommendation models. Our extensive experiments demonstrate that LLM4DSR significantly enhances the performance of both traditional sequential recommendation models and advanced LLM-based models.
The primary contributions of this work are:
Introducing the novel application of LLMs in denoising sequential recommendations and identifying key challenges.
Proposing LLM4DSR, a novel LLM-based recommendation denoising method that utilizes self-supervised fine-tuning and an uncertainty estimation mechanism to fully exploit the capabilities of LLMs.
Conducting comprehensive experiments that demonstrate the superiority of LLM4DSR over state-of-the-art denoising methods across three recommendation backbone models.
2. Preliminary
In this section, we introduce key definitions in this work.
Sequential Recommendation Let be the set of items. We use to denote -th historical temporal interaction sequence where is the -th item interacted within sequence , and represents the sequence’s length. We denote a subsequence of as . For each sequence , sequential recommendation takes sequence data as input and outputs the item that the user is likely to interact with next, which is expected as . For the sake of brevity, we will use to replace in the following text.
Denoise Task Formulation For a given sequence , the objective of the denoising task is to identify and remove the noise within the sequence and ultimately generate a noiseless sequence . Previous explicit denoising tasks simply performed deletion operations on the sequence, resulting in a subsequence that is shorter than the original sequence, i.e., and .
In this paper, the denoising task is accomplished by substituting the identified noise with a more appropriate item. The reason is it can effectively reduce the impact of erroneous deletions, and the extra items introduced by the replacement operation may indeed introduce some useful information benefiting sequence recommendation. We also empirically validate its effectiveness in our experiments. The replacement operation maintains the sequence length unchanged, i.e., and modify the original sequence by
where indicates whether each item at a given position is noise. indicates that the denoising model classify the item as noise, while indicates not. denotes the replacement item generated by the model for .
3. Method
In this section, we outline the methodology of LLM4DSR. Section 3.1 details the training process for the LLM to enable its denoising capabilities. In Section 3.2, we present the uncertainty estimation module, which is designed to enhance the accuracy of the denoising process. Finally, Section 3.3 describes the procedure for replacing noise samples with the suggested items provides by LLMs.
3.1. Training LLM as a Denoiser
Due to the significant discrepancy between the original training objectives in pre-training phase and denoising tasks, LLMs inherently lack the capability for denoising. Previous studies have indicated that fine-tuning with a small amount of data can enable LLMs to quickly adapt to new tasks (Brown et al., 2020). In this section, we design a self-supervised task and employ supervised fine-tuning (SFT) to empower LLMs with denoising ability.
For an original sequence in the training set, we randomly select an element and replace it with an item randomly sampled from the item set , generating a modified sequence . The training objective of the LLM is to identify the replaced item given the modified sequence and predict the original item at that position. This self-supervised task can be expressed in natural language as follows:
Here, represents the title of item . We denote the token sequence of input part as and output part as . The token sequences for ”Noise Items:” and ”Suggested Items:” in the output are defined as and , respectively. Then we can fine-tune the LLMs on the aforementioned corpus and the objective function can be formalized as follows:
where represents the model parameter and is the total length of token sequence of the output part. signifies the probability of the token , given the input sequence and all preceding tokens .
Training on the aforementioned self-supervised tasks empowers the LLMs to effectively understand and perform denoising tasks. The prompt ”The user has interacted with the following items before:” provides the model with information about the user’s historical interactions. ”Noise Items:” guides the model to identify noise within the sequence, as these noise items typically deviate significantly from other items. ”Suggested Items:” encourages the model to infer the user’s interests based on the most items in the sequence, thereby suggesting items that align more closely with the user’s preferences.
While the LLM has developed the capability to denoise using the ”find and replace” task, one limitation is that it can only address a single noise item per sequence. This is because the training data used during the fine-tuning phase contains only one item in the output, leading the model to strictly output one item during the inference phase as well 111We also attempted to perform multiple replacements on a sequence in the self supervised task to enable the LLM to output multiple noise items, but we found that this approach yielded very low accuracy.. Consequently, the model is unable to effectively manage situations where multiple noise items are present. Furthermore, due to the hallucination phenomenon inherent in LLMs, directly relying on the model’s output to identify noise is unreliable. The model may generate fake items that are not actually present in the sequence, or it may lack high confidence in the given item, thereby compromising the accuracy of the denoising results. To mitigate these issues, we have developed an uncertainty estimation module, which calculates the probability of each item in the sequence being noise and use them to identify the noise.
3.2. Uncertainty Estimation
Relying exclusively on responses from LLMs for denoising is both imprecise and inflexible. We developed an uncertainty estimation module designed to gauge the model’s confidence in its response.
For a given sequence , we define the probability of item being noise as . Considering that LLMs are capable of modeling text probabilities, we utilize the LLM to provide the probability of appearing after the prompt ”Noise Items:” in the context of a user-provided sequence . This probability serves as an indicator of the LLM’s confidence in identifying the item as noise and can be treated as an estimation of . Formally,
where denotes the token vector of the item and is the total length of .
The calculation of necessitates a forward pass for each individual item in the sequence, which can be computationally expensive for LLMs. Based on the observation that the first token of different items often varies for most sequences, we can simplify the probability computation. For such sequences, once the first token is given, the subsequent tokens are fully determined. This is because the generated noise item must be included among the sequence items. Formally,
Thus we can get,
Given the aforementioned simplification, we only need to perform a single forward pass using the prompt to obtain the probability values of all items in the sequence being noise, no need for performing one forward pass for each item in the sequence individually.
By modeling text probabilities using a LLM, we are able to determine the likelihood of each item in a sequence being noise. Ideally, a higher probability indicates a greater likelihood of the item being noise. We validated this hypothesis through experiments conducted on a dataset with artificial noise (noise label are available), the details of datasets can be found in Section 4.1.1. As shown in Figure 3, both the number and proportion of noise samples increase as the probability values increase, while the number of clean samples decreases. This experiment confirms the reliability of the probability provided by the LLM.
We propose setting a threshold to classify items with probability values exceeding this threshold as noise. Formally, for a sequence , we denote the probability values for each item as . We introduce a hyperparameter as the threshold and classify items with a probability greater than as noise, i.e.,
Uncertainty estimation of leveraging model-generated probability, rather than relying solely on response-based discrimination, enhances the accuracy of noise detection while also increasing the algorithm’s flexibility. In certain sequences, the model may exhibit high confidence that multiple items are noise, resulting in several items being modified, whereas in others, the confidence levels across all items may be relatively low, leading to no modifications being made.
3.3. Replace Noise with Suggested Items
Instead of simply removing noise, it is more logical to replace it with an item that closely aligns with the user’s preferences. Replacing rather than simply deleting can mitigates the issue of information loss resulting from erroneous deletions. By leveraging the powerful reasoning capabilities and extensive knowledge of LLMs, we can enable the LLM to select an item that better matches the user’s historical interaction patterns to replace the noise sample.
Formally, for a sequence , if is identified as noise (i.e., ), we use the prompt to generate a suggested item, referred to as the oracle. Due to the hallucination phenomenon inherent in LLMs, the oracle may be a fake item that does not necessarily exist in the item set. Inspired by BIGRec (Bao et al., 2023a), we employ grounding techniques to align it with the item set by searching for the item whose embedding is most similar to that of the oracle.
where denotes the embedding of the item and denotes the embedding of the suggested item’s title generated by the LLM. The grounding operation maps the generated item oracle to the actual item with the most similar embedding.
4. Experiments
We aim to answer the following research questions:
RQ1: How does LLM4DSR perforem compare to existing state-of-the-art sequential recommendation denoising methods?
RQ2: What are the impacts of the components (e.g., uncertainty estimation module, replace with suggested items) on LLM4DSR?
RQ3: How does LLM4DSR identify noise and complete the sequence?
4.1. Experimental Settings
4.1.1. Datasets.
Three real-world datasets with 10-core setting: Amazon Video Games, Amazon Toy and Games, and Amazon Movies 222https://jmcauley.ucsd.edu/data/amazon/index_2014.html were utilized in our experiments, which have been used in the previous works (Cui et al., 2024). To ensure a fair comparison, we adopted the data preprocessing used in recent studies (Bao et al., 2023a). For each user sequence longer than 11 interactions, a sliding window of length 11 is used to segment the sequences. The sequences were then organized in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1. We randomly retained 20,000 items for Amazon Movies due to its large size. The dataset statistics are presented in Table 1.
Two noise settings are used in the experiment. The first setting uses the original real-world datasets, which are assumed to inherently contain natural noise. The second setting involves artificial noise, where we randomly replace items in the user interaction sequences according to a specified probability. The noise ratio indicates that each item is randomly replaced with a probability of . Note that the artificial noise is introduced before segmenting the sequences using the sliding window, thus preventing any information leakage between sequences.
| Datasets | #Users | #Items | #Interactions | #Sequences | #Density |
| Game | 54955 | 17256 | 494934 | 149393 | 0.0522% |
| Toy | 19124 | 11758 | 165247 | 47931 | 0.0735% |
| Movie | 9329 | 5101 | 214980 | 121690 | 0.4518% |
4.1.2. Baselines.
We conducted experiments on three backbones, including the traditional sequential recommendation models SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), as well as a LLM-based recommendation model LLaRA (Liao et al., 2024). The methods compared in the study fall into two categories:
Explicit Denoising Methods (HSD (Zhang et al., 2022), STEAM (Lin et al., 2023)): HSD identifies noise based on inconsistency signals, whereas STEAM employs traditional models to train a discriminator and generator for noise identification and completion. 333We acknowledge the existence of other explicit denoising methods, such as SSDRec (Zhang et al., 2024a) and END4Rec (Han et al., 2024). We excluded them from our study because it necessitates additional information, which is incompatible with our experimental setup. Note that we performed end-to-end training on the original backbone (i.e., BERT4Rec) for HSD and STEAM and after completing this training, we transferred the generated denoised dataset to other backbones for training from scratch.
Implicit Denoising Methods (FMLP (Zhou et al., 2022), CL4SRec (Xie et al., 2022)): FMLP introduces FFT to remove noise signals from the representation, and CL4SRec enhances sequences during contrastive learning to increase robustness against noise. We only test implicit denoising methods on the traditional sequential recommendation backbone due to the coupling of implicit denoising methods with specific models, which prevents their application to LLM-based recommendation systems.
4.1.3. Implementation Details.
For training LLM4DSR, we select Llama3-8B (Dubey et al., 2024) as the backbone and randomly sample 5000 prompts to fine-tune the model over 50 epochs using LoRA (Hu et al., 2021). For the threshold hyperparameter , we conduct a search within the [0,1] quantiles of the noise probabilities across all items, with a search interval of 0.05. For traditional models, we fix the learning rate at 0.001 and the embedding size at 64. We use the source code provided in the original papers and search for optimal hyperparameters for all comparison methods according to the instructions in the original papers. The evaluation metrics employed were NDCG@ and Hit Rate@, with set to 20 and 50.
| Games | Toy | Movie | |||||||||||
| Backbone | Method | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 |
| None | 0.0299 | 0.0377 | 0.0646 | 0.1040 | 0.0190 | 0.0239 | 0.0501 | 0.0749 | 0.0468 | 0.0548 | 0.0900 | 0.1308 | |
[…]
| HSD | 0.0246 | 0.0324 | 0.0514 | 0.0908 | 0.0121 | 0.0172 | 0.0319 | 0.0582 | 0.0369 | 0.0451 | 0.0780 | 0.1196 | |
[…]
| FMLP | 0.0369 | 0.0449 | 0.0710 | 0.1116 | 0.0195 | 0.0242 | 0.0426 | 0.0699 | 0.0554 | 0.0641 | 0.0952 | 0.1394 | |
[…]
| BERT4Rec | LLM4DSR | 0.0375 | 0.0464 | 0.0748 | 0.1202 | 0.0222 | 0.0280 | 0.0505 | 0.0799 | 0.0604 | 0.0697 | 0.1046 | 0.1516 |
[…]
| LLaRA | LLM4DSR | 0.0217 | 0.0263 | 0.0352 | 0.0588 | 0.0131 | 0.0169 | 0.0265 | 0.0455 | 0.0166 | 0.0193 | 0.0258 | 0.0392 |
| Backbone | Method | Games | Toy | Movie | |||||||||
| NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | ||
| SASRec | None | 0.0252 | 0.0314 | 0.0526 | 0.0838 | 0.0139 | 0.0178 | 0.0348 | 0.0542 | 0.0390 | 0.0455 | 0.0750 | 0.1084 |
[…]
| HSD | 0.0198 | 0.0250 | 0.0412 | 0.0678 | 0.0090 | 0.0127 | 0.0254 | 0.0444 | 0.0340 | 0.0407 | 0.0700 | 0.1042 | |
[…]
| LLM4DSR | 0.0292 | 0.0357 | 0.0624 | 0.0956 | 0.0152 | 0.0200 | 0.0401 | 0.0642 | 0.0431 | 0.0517 | 0.0876 | 0.1314 | |
| BERT4Rec | None | 0.0266 | 0.0335 | 0.0534 | 0.0880 | 0.0150 | 0.0195 | 0.0346 | 0.0574 | 0.0485 | 0.0569 | 0.0826 | 0.1258 |
| FMLP | 0.0281 | 0.0353 | 0.0598 | 0.0962 | 0.0147 | 0.0188 | 0.0330 | 0.0536 | 0.0478 | 0.0561 | 0.0818 | 0.1240 | |
[…]
| LLM4DSR | 0.0316 | 0.0397 | 0.0644 | 0.1056 | 0.0183 | 0.0231 | 0.0401 | 0.0645 | 0.0509 | 0.0591 | 0.0890 | 0.1306 | |
[…]
| STEAM | 0.0108 | 0.0146 | 0.0224 | 0.0416 | 0.0092 | 0.0120 | 0.0202 | 0.0348 | 0.0107 | 0.0127 | 0.0182 | 0.0286 | |
| LLM4DSR | 0.0150 | 0.0193 | 0.0286 | 0.0504 | 0.0121 | 0.0162 | 0.0280 | 0.0482 | 0.0159 | 0.0186 | 0.0250 | 0.0390 | |
4.2. Performance Comparison (RQ1)
In this section, we analyze the superior of LLM4DSR under two different noise setting as compared with other baselines.
4.2.1. Evaluations on Real-world Datasets.
In Table 2, we present the test results of various methods on real-world datasets. Notably, LLM4DSR consistently achieved the best performance across almost all datasets.
For explicit denoising methods (i.e., STEAM and HSD), we observed that while these methods perform well on BERT4Rec, their effectiveness diminishes when retraining on SASRec using the denoised dataset they generated. For instance, on the Games dataset, the STEAM method improves NDCG@20 from 0.0351 to 0.0373 on BERT4Rec. However, when the same denoised dataset is evaluated on SASRec, NDCG@20 drops from 0.0299 to 0.0233. This suggests that the denoised dataset generated by STEAM lacks good transferability. Additionally, the performance of HSD significantly declined. We argue that this issue may arise from HSD’s inability to effectively capture sequence information in short sequence scenarios. This limitation can result in the loss of critical information when incorrect items are deleted. LLM4DSR also demonstrated impressive performance when applied to LLM-based architectures (i.e., LLaRA). Notably, HSD demonstrates better performance on LLM-based backbones compared to traditional models. This may be due to the fact that the shortened sequences generated by HSD are more compatible with LLMs (Hou et al., 2024), and the inherent reasoning capabilities of LLMs make them relatively insensitive to erroneous deletions.
For implicit denoising methods (i.e., FMLP and CL4SRec), we found that their performance is unstable and can even decline on certain datasets. For example, on the Toy dataset, both FMLP and CL4SRec experienced performance drops, with NDCG@20 decreasing from 0.256 to 0.242 and 0.246, respectively. We hypothesize that for CL4SRec, the substantial difference from the original sequence during the sequence enhancement phase may degrade the quality of positive samples, and using contrastive loss for alignment might introduce additional noise. FMLP, which performs denoising at the representation level through filtering, appears unable to effectively mitigate the impact of noise samples within the sequence.
4.2.2. Evaluations on Artificially Noisy Dataset.
The test results for various methods on the artificially noise-added dataset, with a noise ratio of 10%, are presented in Table 3. These results indicate that LLM4DSR consistently outperforms all comparative methods. Artificial noise, being more distinguishable than natural noise in real-world datasets, makes the improvement of LLM4DSR more pronounced on such datasets. For example, on the Game dataset, LLM4DSR combined with SASRec achieves a 5.3% improvement in NDCG@20, whereas on the artificially noise-added dataset, the improvement reaches 15.8%. Furthermore, we evaluated the performance of all comparative methods under different noise ratios, as illustrated in Figure 4. LLM4DSR significantly surpasses other comparative methods demonstrate its robust performance across varying noise conditions.
4.2.3. Evaluation of the accuracy of explicit denoising methods in identifying artificial noise.
The explicit denoising method achieves denoising by identifying and removing noise, with the accuracy of noise identification being closely related to its overall effectiveness. To better understand the performance of explicit denoising methods, we evaluated the F1 score of various explicit denoising methods on datasets with artificially introduced noise, where noise labels are available. As illustrated in Figure 5, we compare LLM4DSR with two other state-of-the-art explicit denoising methods on the Game and Toy datasets with different noise ratios of 10%, 20%, and 30%. Among these methods, LLM4DSR consistently achieved the highest performance, suggesting that leveraging the open knowledge of LLM can significantly enhance noise identification.
4.3. Ablation Study (RQ2)
To verify the contributions of various components of LLM4DSR, we conducted an ablation study on different variants under SASRec, using both real and artificially noisy datasets from the Game and Toy datasets. The variants tested include: (1) ”w/o uncertainty,” which omits uncertainty estimation and directly uses the LLM output for denoising; (2) ”w/o completion,” which excludes the completion capability of LLM4DSR and performs only deletion operations on sequences; and (3) ”w/ popularity completion,” which replaces noise with items randomly sampled from the item set based on popularity weights. As presented in Table 4, the consistently superior performance of LLM4DSR across all datasets highlights the effectiveness of each component. In particular, the performance decline observed in the ”w/o completion” and ”w/ popularity completion” emphasize the critical role of the replacement operation, demonstrating that the LLM4DSR can effectively extract user interests and provide items that better align with the sequence.
| Dataset | Method | NDCG@20 | HR@20 |
| Game | SASRec | 0.0299 | 0.0646 |
| w/o uncertainty | 0.0309 | 0.0648 | |
| w/o completion | 0.0241 | 0.0566 | |
| w/ popularity completion | 0.0239 | 0.0460 | |
| LLM4DSR | 0.0315 | 0.0696 | |
| Toy | SASRec | 0.0190 | 0.0501 |
| w/o uncertainty | 0.0191 | 0.0501 | |
| w/o completion | 0.0185 | 0.0478 | |
| w/ popularity completion | 0.0156 | 0.0417 | |
| LLM4DSR | 0.0208 | 0.0557 | |
| Game w/ artificial noise | SASRec | 0.0252 | 0.0526 |
| w/o uncertainty | 0.0277 | 0.0572 | |
| w/o completion | 0.0264 | 0.0548 | |
| w/ popularity completion | 0.0241 | 0.0472 | |
| LLM4DSR | 0.0292 | 0.0624 | |
| Toy w/ artificial noise | SASRec | 0.0139 | 0.0348 |
| w/o uncertainty | 0.0133 | 0.0361 | |
| w/o completion | 0.0142 | 0.0359 | |
| w/ popularity completion | 0.0106 | 0.0267 | |
| LLM4DSR | 0.0152 | 0.0401 |
4.4. Case Study (RQ3)
In this case study, we illustrate how LLM4DSR leverages open knowledge to achieve denoising. As depicted in Figure 6, we analyze a user’s viewing history from the Movie dataset, which predominantly consists of films in the drama, romance, and comedy genres. A notable outlier in this collection is the science fiction film ”Escape from the Planet of the Apes”. Utilizing prior knowledge about movie classifications, the LLM identifies this film as noise with a probability of . To enhance the sequence information, LLM4DSR recommends ”The Man from Snowy River”, a film that aligns with the user’s preferred drama/romance genres, as a suitable replacement for the identified noise.
5. Related Work
5.1. Sequential Recommendation
The sequential recommendation infers the next item of interest to the user based on their historical interactions. Compared to collaborative filtering, sequential recommendation takes into account the temporal order of interactions to better capture the dynamic changes in user interests. Benefiting from the development of deep learning models in recent years, sequential recommendation employs various deep models to model users’ historical interactions and capture their interests. For instance, GRU4Rec (Hidasi et al., 2015) uses RNN, while Caser (Tang and Wang, 2018) uses CNN. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are based on the self-attention mechanism (Vaswani, 2017), which automatically learns the weight of each interaction. DROS (Yang et al., 2023) introduces distributionally robust optimization (DRO) (Rahimian and Mehrotra, 2019) to enhance the model’s ability to handle out-of-distribution (OOD) problems. The readers may refer to the excellent survey (Fang et al., 2020; Wang et al., 2019) for more details. However, the aforementioned methods do not consider the negative impact of noisy interactions on modeling the sequence, which is quite common in real-world scenarios.
5.2. Denoising Recommendation
The primary objective of denoising sequential recommendation is to enhance the performance of recommendation systems by mitigating the impact of noisy data. Denoising methods for sequential recommendation can be broadly categorized into implicit denoising and explicit denoising. Implicit denoising methods primarily reduce the impact of noise without explicitly removing them. Some works reduce the influence of noisy interactions to the sequence representation by increasing the sparsity of attention weights in attention-based models (Yuan et al., 2021; Chen et al., 2022). FMLP and END4Rec (Zhou et al., 2022; Han et al., 2024) consider using filters to remove noise signals from representations. The drawback of these methods is that noise still affect the model, and their high coupling with the model architecture makes it difficult to transfer them to other backbones or downstream tasks, including the widely researched LLM-based recommendation systems in recent years.
Explicit denoising methods directly modify the sequence content, effectively addressing the limitations of implicit denoising techniques. These methods often rely on heuristically designed denoising modules or metrics. BERD (Sun et al., 2021) models the uncertainty of each interaction and eliminates modules with high loss but low uncertainty to achieve denoising. HSD (Zhang et al., 2022) assesses the similarity of each sample to other items in the sequence and its alignment with user interests to detect noise. STEAM (Lin et al., 2023) trains a discriminator model and a generator model, which are used to identify noise and complete sequences, respectively.
Regardless of whether explicit or implicit denoising methods are used, the most significant challenge stems from the lack of effective supervision signals from noise labels to train the denoising model. Existing methods often rely on heuristic designs inspired by prior knowledge to develop denoising algorithms, which requires substantial expert knowledge and lacks precision. Additionally, since these methods usually utilize supervision information from recommendation loss, they are still adversely affected by noise as the labels of the recommendation label (i.e., the next item) could also be contaminated. To address these issues, LLM4DSR leverages the extensive open knowledge of Large Language Models (LLMs) to aid in more effective denoising.
Additionally, some works focus on denoising under collaborative filtering scenarios (Wang et al., 2021a, c; Gao et al., 2022; Tian et al., 2022; Wang et al., 2022; Fan et al., 2023; Wang et al., 2023a; He et al., 2024; Li et al., 2024c). Compared to denoising in sequential recommendation, these methods do not take the temporal information of interactions into account. Readers can refer to the survey (Zhang et al., 2023a) for more details.
5.3. LLMs for Recommendation
Large Language Models (LLMs), with their extensive knowledge, generalization capabilities, and reasoning abilities, have achieved remarkable results across numerous fields. LLMs are utilized in various recommendation tasks, including collaborative filtering (Zhu et al., 2024; Wang et al., 2024c), sequential recommendation (Hou et al., 2023, 2022; Bao et al., 2023a), graph-based recommendation (Wei et al., 2024), and CTR (Click-Through Rate) tasks (Bao et al., 2023b; Wang et al., 2023b; Xu et al., 2024). Additionally, LLMs have been utilized for tasks that were challenging for traditional model to achieve, such as explaining item embeddings (Tennenholtz et al., 2023), explaining recommendation results (Gao et al., 2023; Wang et al., 2024d), and conversational recommendations (Gao et al., 2023).
The initial attempts directly use pre-trained LLMs as the backbone for recommendations to explore their zero-shot capabilities. (Gao et al., 2023; Hou et al., 2024; Liu et al., 2024; Wang and Lim, 2023; Wang et al., 2024b). These methods perform not well due to the significant gap between recommendation tasks and the training tasks of LLMs. Subsequent research has organized recommendation data into prompt formats and then fine-tuned LLMs to enhance their recommendation performance, achieving better results (Bao et al., 2023a, b; Hu et al., 2024; Li et al., 2023a, 2024b, b, 2024a; Shi et al., 2024; Wang et al., 2024a; Zhang et al., 2023b, 2024b). Besides generating recommendation results directly, LLMs can also serve as enhancers to improve the performance of traditional recommendation models (Xi et al., 2023; Ren et al., 2024; Wei et al., 2024; Yang et al., 2024; Wang et al., 2023b).
To the best of our knowledge, this paper is the first to explore the application of LLMs in the field of sequence recommendation denoising.
6. Conclusion
In this paper, we propose LLM4DSR, a novel method for sequence recommendation denoising using Large Language Models (LLMs), which is the first exploration of LLM performance on this task. Leveraging the extensive open knowledge embedded in LLMs, LLM4DSR offers significant advantages over traditional denoising methods. We have designed a self-supervised task that endows LLMs with enhanced denoising capabilities and integrated the Uncertainty Estimation module to improve both the accuracy and flexibility of the denoising process. Furthermore, we utilized the generative capabilities of LLMs to complete the denoised sequences, thereby enriching sequence information and ensuring robust performance in downstream tasks. Comprehensive experiments demonstrate that LLM4DSR outperforms existing state-of-the-art sequence denoising techniques.
Despite the notable performance of LLM4DSR, it incurs higher computational costs compared to traditional methods due to the use of LLMs. A promising direction for future research is to employ techniques such as knowledge distillation to transfer the denoising capabilities of LLMs to smaller models, thereby achieving superior performance without increasing computational overhead.
Ethical Considerations
Denoising methods essentially transfer the model’s prior knowledge to the data. If the model’s outputs are biased, for example, if it tends to recommend items of a specific category, it will further amplify the biases present in the data. In the context of using LLMs for denoising as discussed in this paper, it is crucial to ensure that the outputs of the pre-trained LLM are inherently fair. Special attention must be paid to bias testing when selecting the backbone model.